A package of 6 fuses are tested where the probability an individual fuse is defective is
0.05. (That is, 5% of all fuses manufactured are defective). What is the probability
that more than one fuse will be defective, given that at least one is defective?
1
Expert's answer
2020-06-08T19:50:34-0400

"P(X>1|X\ge1)=\frac {P(X>1) \bigcap P(X\ge 1)}{P(X\ge 1)}"

="\frac {P(X>1)}{P(X\ge 1)}"

For a binomial probability,

"P(X=x) ={n\choose x} *p^x *q^{n-x}"

P(X>1)="\sum_{i=2}^6{6\choose x} *0.05^x *0.95^{6-x}"

=0.0328

"P(X\ge 1)=\sum_{i=1}^6{6\choose x} *0.05^x *0.95^{6-x}"

=0.2649

Then

"\frac {0.0328}{0.2649}"

=0.1238